Session: May 1998 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Alessandro Cattani at Mon Jun 1 12:23:13 METDST 1998) Success
We have terminated the experiment with success

Jodrell (Alastair Gunn at Mon Jun 1 15:45:36 METDST 1998) Minor failures
Receiver cryostat system failed prior to this experiment. Gain was much reduced but weak fringes were seen by MERLIN. Telescope off-line between 08:03 UT and 09:21 UT. Had some tape head position errors but values look okay.

Cambridge (Alastair Gunn at Mon Jun 1 15:46:07 METDST 1998) Minor failures
Had vacuum problems at start due to rc communications errors. FS re-started at 07:40 UT. Missed first four scans, started okay at 07:50 scan. No other known problems.

Onsala (A. Polatidis at Mon Jun 1 21:05:46 METDST 1998) Failure
Onsala did not participate in this epxeriment due to failure of the declination brake in the 25m antena. The problem was investigated today and we will be observing the next experiment.

Torun (K. Borkowski at Tue Jun 2 09:22:56 METDST 1998) Success
No known problems.

Westerbork (Tony Foley at Tue Jun 2 10:48:25 METDST 1998) Minor failures
Late onsource during the first scan. Decoder still giving problems.

Effelsberg (dgraham at Tue Jun 2 18:45:15 METDST 1998) Minor failures
good enough

Urumqi (Liu Xiang at Wed Jun 3 12:25:58 METDST 1998) Minor failures
The IF total power is not high enough, i.e. about 5000-10000k.

Calibration (Jean\-Francois Desmurs at Mon Jun 8 13:30:08 METDST 1998) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name c98c2.antab. Rem : Ur: Bad Tsys values. Read the comments in the file to get further information.

Noto (C.Trigilio at Mon Jun 8 17:33:30 METDST 1998) Minor failures
The Formatter has been found one sec late; it is not clear if the problem occurred during the observation or if it was set wrong (with fmset) before the start, during the FS reconfiguration from VLBA to MK3. Some problem with BBC03, that sometimes was found with no_1pps.

Urumqi (Liu Xiang at Wed Jun 10 10:50:59 METDST 1998) Minor failures
The Phase-Call signal too strong, so we switched off it. The Noise Injet failed, we switched off it. The Clock Offset is UTC: 06:30---1.9us, UTC: 07:05---CLOCK SIGNAL JUMP, UTC: 07:20---Clock Offset is 2.1us. The tape was sent in 3, Jun.